a cannon of mass 5m ( including a shell of mass m )is at rest on a smooth horizontal ground fires the shell with its barrel at an angle A with the horiZontal at a velocity u relative to the cannon. Find the horiZontal distance of the point where shell strikes the ground from the initial position of the cannon -
I need the procedure

Grade:12

2 Answers

tejas naresh patel
10 years ago

345

Akash Kumar Dutta
10 years ago

since initial momentum is 0 final must also be,
relative=u
hence absolute=u + v...where v=recoil velocity,
 4m.v=m(u+v)
 v=u/3
hence initial velocity=u+u/3=4u/3
and angle of projection=A
hence Range=  16 u^2 sin2A
                        ____________    ANS
                             9g